OpenStudy (anonymous):

Part a.) Write the equation of the line that has a slope of -2 and a y-intercept of -6. (5 points) Part b.) Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3. (5 points)

OpenStudy (anonymous):

part a) y-intercept slope "y=mx+b" y=mx+b m=slope b=y-intercept slope -2 y-intercept -6 so you put your numbers into the equation y=-2x-6 and that's your answer part b) you need point slpe form which is y-y1=m(x-x1) m= slope, and you have (6,-2) which is your (x1,y1) put them into your equation y-(-2)=-2/3(x-6) or y+2=-2/3(x-6) multiply -2 on the left and the 3 to rigth y-4=3x-18 add 4 both sides y=3x-14 that's you answer